CAVITY RING-DOWN LASER ABSORPTION SPECTROSCOPY OF YTTRIUM AND NICKEL MONOIO-DIDES

TONGMEI MA, LING LI, J. W-H. LEUNG, A. M-Y. WONG and \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.